Multiple choice. 18,19,20, and 21

Multiple Choice Identily the choice that best completes the statement or answers the question. 18. What property displays text onscreen for a label control? a. Name b. ControlSource c. TextAlign d. Caption 19. Study the following function that calculates net pay. Function NetPay(GrossPay,FederalTaxes, StateTaxes, LocalTaxes) NetPay-GrossPay-(FederalTaxes StateTaxes LocalTaxes) End Function Given the values for GrossPay#800. FederalTaxes#75, StateTaxes-20, LocalTaxes 10, what is the value returned by the function? a 855 b. 695 С. 905 d. 705 20. which of these describes a istbox? a, indicates whether an option isc. Displays a list of values from which the user can select one selected or not b. accepts data from the user d. displays text on the user form 21 Which one is true about a Dynamic Array? a The Array size is unknown b. The Array size is 2000 c The Array size is 1200 d. The Array size is 1000
Answer #1

18)

ControlSource is used to display onscreen for a label control

19)

netpay = 800 - (75 + 20 + 10) = 800 - 105 = 695

Answer: option(b)

20)

listbox allows users to select a box from items in list

which conatins multiple text boxes.

Answer: option(c)

21)

Dynamic array as goes by the name is dynamic in nature with no defined size

hence size is unknown

Answer: option(b)
Answer #2
The answer to 19 is B.
answered by: anonymous